Two round rods , one of steel , the other of brass, are joined end to end. Each rod is 0.5 m long and 2 cm in diameter. The combination is subjected to a tensile force of 5000 N. Find the strain in each rod.

    Young’s Modulus of Steel use 20 x 10^10 Pa,of brass use 10 x 10^10 Pa.


1
Expert's answer
2022-05-02T13:43:26-0400

"s_i=\frac{1}{\Upsilon_i}\frac {4F_i}{\pi d_i^2},"

"s_1=8\cdot 10^{-5},"

"s_2=4\cdot 10^{-5}."